The length was 8 ft. more than twice the width. If the perimeter was 136ft. find the dimensions.
For the equation i got 8(2x)+8(x)
Is this right if not can you help me

Answer by ankor@dixie-net.com(22740)

The length was 8 ft. more than twice the width. If the perimeter was 136ft.
find the dimensions.
:
Not sure of what you have done here, but this is the way I would do it:
:
The perimeter
2L + 2W = 136
:
"The length was 8 ft. more than twice the width.", therefore
L = 2W + 8
:
In the perimeter equation, replace L with (2W+8)
2(2W+8) + 2W = 136
4W + 16 + 2W = 136
4W + 2W = 136 - 16
6W = 120
W = 120/6
W = 20
Find L
L = 2(20) + 8
L = 48
:
:
Check this in the perimeter equation